Bolander Corporation uses a job-order costing system with a single plantwide predetermined overhead rate based on machine-houThe following data have been recorded for recently completed Job 450 on its job cost sheet. Direct materials cost was $3,044.
Answer #1
1
Predetermined overhead rate 6.50 =2.3+(294000/70000)
Direct materials 665
Direct labor cost 1840
Overhead applied 520 =80*6.50
Total job cost for Job M825 3025
Option C $3,025 is correct
2
Direct materials 3044
Direct labor cost 690 =46*15
Overhead applied 1352 =104*13
Total job cost for Job 5086
Option B $5,086 is correct